OpenStudy (anonymous):

how do i simplify the square root of 84

OpenStudy (cwrw238):

the square root of a number N is a number which when multiplied by itself gives N. if you find the prime factors of 81 the square root will be apparent: 81 = 3*3*3*3

OpenStudy (anonymous):

i meant 84

OpenStudy (anonymous):

but i think the answer is 2 sq root of 42 that is one of the answer choices

OpenStudy (cwrw238):

oh ok well factor 84 and look for perfect squares 84 = 2*2*3*7 = 4* 21 squareroot of 4 = 2 answer is 2 sqrt21

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

you simplify it by simplifying